OpenStudy (anonymous):

find the equation of the circle that has the points P(-3,1) and Q(1,3) as endpoints of a diameter

OpenStudy (anonymous):

midpoint of these two will be the center. distance from the center to one of these points will be the radius. you can then plug all that into the form y=(x-h)^2+(y-k)^2=r^2 where (h,k) is the center and r is the radius